OIL PRICE SHOCKS AND LONG RUN PRICE AND IMPORT DEMAND BEHAVIOR

FRANK KLEIBERGEN^{1}, HERMAN K. VAN DIJK^{2} AND JEAN-PIERRE URBAIN^{3}

1Econometric Institute, Erasmus University Rotterdam, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands 2Econometric and Tinbergen Institutes, Rotterdam, Burg. Oudlaan 50, 3062 PA Rotterdam, The Netherlands 3Department of Quantitative Economics, University of Maastricht, P.O. Box 616, 6200 MD Maastricht, The Netherlands

(Received April 5, 1995; revised November 26, 1997)

Abstract. The effect which the oil price time series has on the long run properties of Vector AutoRegressive (VAR) models for price levels and import demand is investigated. As the oil price variable is assumed to be weakly exogenous for the long run parameters, a cointegration testing procedure allowing for weakly exogenous variables is developed using a LU decomposition of the long run multiplier matrix. The likelihood based cointegration test statistics, Wald, Likelihood Ratio and Lagrange Multiplier, are constructed and their limiting distributions derived. Using these tests, we find that incorporating the oil price in a model for the domestic or import price level of seven industrialized countries decreases the long run memory of the inflation rate. Second, we find that the results for import demand can be classified with respect to the oil importing or exporting status of the specific country. The result for Japan is typical as its import price is not influenced by GNP in the long run, which is the case for all other countries.

Key words and phrases: Cointegration, weak exogeneity, import demand, oil price behaviour.